You plan to open a Subway restaurant.  The franchise fee is $15,000 and startup costs are an additional $150,000.  The average customer spends $18.  Variable costs (food, labor, etc.) are $14 per customer. How many customers do you need to serve to break even?
Answer #1

VALUES:
FIXED COST = 165000
VARIABLE COST = 14
PRICE PER UNIT = 18

BREAK-EVEN QUANTITY = FIXED COST / CONTRIBUTION MARGIN

CONTRIBUTION MARGIN = PRICE PER UNIT - VARIABLE COST = 18 - 14 = 4

BEQ = 165000 / 4 = 41250